OpenStudy (anonymous):

Solve t divided by fifteen equals one and two-fifths. seven seventy-fifths three and one-fifth 3 twenty-one

OpenStudy (anonymous):

The multiplication rule for fractions is quite easy to remember: Multiply numerators and then multiply the denominators. [a over b] times [c over d] example where b ≠ 0, d ≠ 0 The development of the rule for multiplying a over b and c over d actually requires two steps: multiplying c over d by a and dividing c over d by b. Multiplication by a counting number involves repeated addition. Example: General rule The general rule is a × c over d = a times c over d where d ≠ 0. Division of a fraction by a counting number takes more thought. What is two-fifths ÷ 3? Think of two-fifths as two unit fractions of one-fifth. Three does not divide into 2 evenly. However, we can write two-fifths as [2 × three-fifths × 3] = six-fifteenths. This is 6 unit fractions of one-fifteenth, and 3 one-fifteenth divided by 3. So the quotient equals 2 unit fractions of one-fifteenth, or two-fifteenths. The general rule is c over d ÷ b = c over b × d where b ≠ 0, d ≠ 0. Notice that when d = 1, this shows that c ÷ b = c over b. Using what has been shown above, a over b × c over d = a × c over b × d can be derived. First divide the fraction c over d by b and then multiply by a. c over d ÷ b = c over b × a and a × c over b × d = a × c over b × d.

OpenStudy (anonymous):

sooo

OpenStudy (anonymous):

@bigrob30117

OpenStudy (anonymous):

its D 21